OpenStudy (anonymous):

which is a factor of 100a b squared - 4 a squared answers : a)a+3 b)a-3 c)a-10 d)a+5 e)5b+a

OpenStudy (anonymous):

can you write the equation using the equation maker. Cannot understand it.

OpenStudy (unklerhaukus):

\[100ab^2-4a^2\]?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

well you have two terms, factorise the both, \[100ab^2=4\times25\times a\times b^2=2^25^2ab^2\] \[4a^2=2\times2\times a^2=2^2a^2\]

OpenStudy (anonymous):

\[4(25b ^{2}-a ^{2})\] \[4((5b)^{2}-a ^{2})\] Now Breaking using formula a^{2}-b^{2} we get \[4(5b+a)(5b-a)\] Therefore answer is 5b+a